OpenStudy (anonymous):

Find two numbers such that their sum is 24 and their product is 143?

OpenStudy (anonymous):

x + y = 24 x = 24 - y y (24 - y) = 143 24y - y^2 = 143 y^2 - 24y + 143 = 0 Solve for y

OpenStudy (mertsj):

(y-11)(y-13)=0 y-11=0 or y-13 = 0 y = 11 or y = 13

OpenStudy (anonymous):

\[x+1/4 +3x+2/2=5\]

OpenStudy (mertsj):

4x+5/4=5 4x=5-5/4 4x = 15/4 x = 15/16